OpenStudy (anonymous):

Solve the Equations Beloww ; 2/5n+6=10 -7-6y=13

OpenStudy (mr.math):

Is it \(\frac{2}{5}n+6=10\)?

OpenStudy (anonymous):

Yeahh '

OpenStudy (mr.math):

1- Subtract 6 from both sides: \[\frac{2}{5}n=4\] 2-Multiply both sides by \(\frac{5}{2}\): \[n=10.\]

OpenStudy (mr.math):

Try the second one.

OpenStudy (anonymous):

Thankss